OpenStudy (anonymous):

Evaluate: the absolute value of the quantity of 2 point 8 minus k + 3 point 4 if k = 8.1 A. 8.7 B. -1.9 C. 14.3 D. 1.9

OpenStudy (anonymous):

2.8-k+3.4 = 2.8-8.1+3.4= -5.3+3.4= -1.9

OpenStudy (anonymous):

B

OpenStudy (anonymous):

wait the absolute value of that is 1.9